OpenStudy (anonymous):

Find f(x) if:

OpenStudy (anonymous):

\[f \prime (x) = 7/ x^4 and f (1) = 4. \]

OpenStudy (anonymous):

f′(x)=7/x4 and f(1)=4 ?

OpenStudy (anonymous):

You just need to integrate f'(x), without forgetting the constant of integration, and then put in the know values to find it, giving you the complete f(x)

OpenStudy (anonymous):

\[\int\limits_{}^{}7/x^4dx = 7\int\limits_{}^{}1/x^4dx = 7\int\limits_{}^{}x^-4dx\]

OpenStudy (anonymous):

that last one is x to the negative fourth. If you can integrate that, you should be fine